a random sample of 429 college students was interviewed about a number of matters. use the results to construct confidence interval estimates of the population mean at the 99% level.

A. they reported that they had sent an average of $478.23 on textbooks during the previous semester , with a sample standard deviation of $15. 78

B. They also reported that they had visited the health clinic an average of 1.5 times a semester with a sample standard deviation of 0.3

C. on the average, the sample had missed 2.8 days of classes per semester because of illness with a sample standard deviation of 1.0

D. on the average, the sample had missed 3.5 days of classes per semester for reason other then illness, with a sample standard deviation of 1.5

please use excel


1
Expert's answer
2022-03-14T07:23:39-0400

N=429, 99% confidence interval for the population mean is "μ±Z∙s/\sqrt{(N-1)}"

The critical value at 0.01 is 2.58

A. "\mu=478.23,s=15.78"

confidence interval = "478.23±2.58(15.78/\sqrt{(429-1)})"

"=478.23\u00b11.97"

"=[476.26,480.2]"

Therefore the student spends between $476.26 and $480.20 on books.


B. "\mu=1.5,s=0.3"

confidence interval "=1.5±2.58(0.3/\sqrt{(429-1)})=1.5±0.04"

"=[1.46,1.54]"

Therefore the students visited the clinic between 1.46 and 1.54 times on average.


C. "\mu=2.8,s=1.0"

Confidence interval "=2.8±2.58(1.0/\sqrt{(429-1)})=2.8±0.12"

"=[2.67,2.93]"

Therefore the students missed classes per semester due to illness on the average between 2.67 and 2.93 times.


D. "\mu=3.5,s=1.5"

"=3.5±2.58(1.5/\sqrt{(429-1)})=3.5±0.19"

"=[3.31,3.69]"

Therefore the students missed classes per semester due to other reasons than illness on an average between 3.31 and 3.69.